# Imaging nuclei by smashing them at high energies: how are their shapes revealed after destruction?

### v2

@@ Abstract
High-energy nuclear collisions has recently emerged as a powerful “imaging-by-smashing” tool to reveal the global shapes of atomic nuclei. Here, I layout a conceptual framework for this technique, explaining how nuclear shapes are encoded during quark-gluon plasma formation and evolution, and how they can be decoded from final-state particle distributions. I highlight the method’s potential to advance our understanding of both nuclear structure and quark-gluon plasma physics.
@@
Imaging techniques are fundamental to scientific discovery, enabling us to visualize and understand structures across vastly different scales. These methods employ external probes, often photons, to interact with samples and glean structural information. The sample’s structure information is encoded in the scattered probe, such as in X-ray crystallography, where a molecule’s three-dimensional organization is deduced from momentum-space scattering patterns. Imaging process takes place as the probe traverses the sample, and because any structural disintegration happens only after scattering, the captured information remains intact.
@@
A radially different approach has emerged in nuclear physics: rather than using external probes, atomic nuclei are collided at ultra-high energies to create a hot, dense state of matter called the quark-gluon plasma (QGP). Remarkably, the geometric structure of the colliding nuclei determines the QGP’s initial conditions and influences its subsequent evolution, ultimately imprinting specific patterns on the momenta of thousands of final-state particles. By carefully “rewinding” this evolution, one can reconstruct an effective image of the original nuclei.
@@
The STAR Collaboration demonstrated this concept by successfully extracting the shape of 238U from 238U+238U collision data. Despite completely destroying the nuclei, their structural information is preserved with sufficient detail to allow reconstruction STAR Collaboration (2024). This breakthrough raises fundamental questions: How are nuclear shapes encoded in such violent collisions? how can they be deduced from final-state particles? What are the broader implications for nuclear physics?
@@
Traditional methods for nuclear shapes. Atomic nuclei are bound systems of protons and neutrons (nucleons) held together by the strong nuclear force. While often depicted as spheres, nuclei can adopt prolate, oblate, or even pear-like configurations – shapes that reflect their underlying many-body wavefunctions. These wavefunctions governs nucleon positions and momenta. The non-perturbative nature of strong force makes theoretical predictions of these shapes challenging and experimental input crucial.
@@
Unlike molecules, nuclei cannot be oriented in a crystalline lattice for direct coherent imaging. Instead, low-energy techniques such as electron scattering, laser spectroscopy, and Coulomb excitation (Coulex) have been used to infer nuclear shapes Yang et al. (2023). Electron scattering probes nuclei one at a time, providing an orientation-averaged image where nuclear deformation appears as modifications to charge distributions. Laser spectroscopy probes nuclear shapes through measurements of hyperfine atomic transitions. Coulex uses low-energy ions to excite nuclei into rotational and vibrational states, where nuclear shapes are deduced by detecting gamma rays emitted during relaxation to ground state and comparing the data to theoretical models.
@@
However, these methods rely on electromagnetic interactions and thus cannot directly probe neutron distributions. Moreover, their relatively long timescales yield time-averaged, blurred representations of nuclear shapes, akin to long-exposure photographs. This temporal averaging obscures rapid fluctuations and dynamic variations in nuclear shapes.
@@
Initially, the nucleus’s shape directly influences the conditions under which the QGP forms. As the plasma expands hydrodynamically, information about the original nuclear shape transforms and becomes encoded in the momentum distribution of thousands of final-state particles. By analyzing these distributions and reverse-engineering the expansion process event-by-event, nuclear shapes can be reconstructed. The precision of this method hinges on how well the initial condition, equation of state, and transport properties of the QGP are understood. These aspects have been the focus of the heavy-ion community for decades and are now believed to be well-constrained Busza et al. (2018).
@@
Despite its seemingly extreme nature, high-energy nuclear imaging offers distinct advantages. First, the collision probes the ground state nuclear many-body wavefunction in position space, information not readily accessible through more conventional means. The colliding nuclei and their Coulomb fields are Lorentz-contracted, ensuring that the nuclei are causally disconnected before the collision. This causal disconnection prevents pre-collision excitation effects that could contaminate the measurement. While nuclei could be Coulomb excited when they pass by each other, most isomeric states are very short-lived (typically $<10^{-12}$ s) and cannot persist in the circulating beam long enough to participate in the subsequent collisions  [^1] Some isomers have exceptionally long lifetime, such as 7.7s 11/2- state of 197Au and 8.9 days 11/2- state of 129Xe, which can accumulate in the beam. However, no estimate are currently avaliable on their production rates in high-energy accelarators.
@@
Energy dependence of the nuclear image. The observed nuclear image, and its effective many-body wavefunction, depends critically on the collision energy $\sqrt{s_{\mathrm{NN}}}$, which determines the effective “shutter speed” of the imaging process. At low energies, where the shutter speed is slow, collective rotational and vibrational DOF dominate the nucleus’s apparent shape. For instance, the rotational DOF of a deformed heavy nucleus, with excitation energies around 0.1 MeV, corresponds to timescales of $\tau\sim 10^{3}-10^{4}$ fm/$c$. As collision energy increases, faster modes such as nucleon clustering and short-range correlations come into play. At still higher energies, the resonance structure of nucleons emerges, eventually giving way to the subnucleonic quarks and gluons. The imaging process effectively captures all DOF slower than the nuclear crossing timescale. As a result, varying $\sqrt{s_{\mathrm{NN}}}$ provides a natural way to study the evolution of the nuclear wavefunction across energy scales
@@
While electron-nucleus scattering probes the one-body distribution of the scattering centers – protons at low energy and quarks and gluons at high energy – much less is known about the nuclear many-body distributions across energy scales. High-energy heavy-ion collisions help fill this gap. Measurements of anisotropic flow coefficients reveal quantum fluctuations associated with the finite number of scattering centers Heinz (2013). This quantum fluctuations induces deformation in transverse ($xy$) nucleon distribution, with eccentricities scaling roughly as $1/\sqrt{A}$, where $A$ is the mass number  [^2] The eccentricities are mostly determined by the positions of nucleons, as most quarks and gluons are confined within them and nuclear modifications are modest.. This fluctuation-driven deformation produces non-zero flow even in head-on collisions Chatrchyan et al. (2014), despite nuclear overlap region being isotropic on average (see Fig. 2).
@@
Figure 3: Relation between initial- and final-state of high-energy nuclear collisions. The features of collision geometry are characterized by its shape and size parameters, $\mathcal{E}_{n}$ and $d_{\perp}$ (${\bf a}$). They are linearly related to observables that describe the transverse momentum $p_{\mathrm{T}}$ spectra in each event, i.e., the anisotropic flow coefficients $V_{n}$ and the average transverse momentum $[p_{\mathrm{T}}]$ (${\bf b}$). The event-to-event variations of these initial- and final-state quantities are linearly related.
Reverse-engineering nuclear shape. A nucleus with quadrupole deformation can be described by a surface function in terms of the polar angle $\theta$ and azimuthal angle $\phi$,
@@
Figure 4: Three steps in the imaging-by-smashing method. llustrated here are collisions of spherical nuclei (${\bf a1}$${\bf a3}$) and prolate-deformed nuclei with $\beta_{2}=0.28$ (${\bf b1}$${\bf b3}$) for two representative collision events: (${\bf a1}$ and ${\bf b1}$) initial configurations of the colliding nuclei, (${\bf a2}$ and ${\bf b2}$) initial geometry in the transeverse plane, and (${\bf a3}$ and ${\bf b3}$) final-state distribution of particles in azimuthal angle and $p_{\mathrm{T}}$. The event-to-event variation in the initial-state geometry and final-state distributions are quantified by the moments involving $\varepsilon_{n}$ and $d_{\perp}$ of the initial state and $v_{n}$ and $[p_{\mathrm{T}}]$ of the final state, as indicated in the middle row. Parameter $a$ represents the nuclear surface diffusivity. The nucleon positions are simulated in a Monte-Carlo Glauber model with $R_{0}=6.81$ fm, $a=0.55$ fm and $A=238$. Imaging involves reconstructing the initial geometry from final-state particles (Step 1), relating it to the nuclear shape, which is affected by both global deformation and quantum fluctuations (Step-2), and comparing different collision systems to isolate the global deformation (Step-3).
[…]
@@
[…]
Event-by-event variations in nuclear geometry, arising from both global deformation and quantum fluctuations, are quantified by moments of initial- and final-state observables. These moments are connected by the linear response relations shown in Fig. 3. Our understanding of the initial conditions, the evolution and properties of the QGP are mainly derived from the study of these relations Heinz (2013); Busza et al. (2018). Many such moments can be used, as indicated in the middle row of Fig. 4. The STAR collaboration specifically focused on three: $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ STAR Collaboration (2024).
@@
Step2: from the initial condition to nuclear shape. The QGP’s initial condition is closely tied to nucleon distributions in colliding nuclei A and B, described by thickness functions $T_{\rm A,B}(x,y)$, which fluctuate from event to event. In the presence of global deformation, these functions also depend on the nuclei’s orientations prior to collision (see Fig. 4a1 and 4b1). For head-on collisions, random orientations of prolate deformed nuclei lead to significant, anti-correlated fluctuations between $v_{2}$ and $[p_{\mathrm{T}}]$ Giacalone (2020). General considerations imply that such fluctuations follow simple parametric forms Jia (2022b). Specifically, the three moments used by the STAR Collaboration are
@@
Here $a_{n}$ and $b_{n}$ are positive coefficients that depend on the impact parameter. The $a_{n}$ terms capture contributions from quantum fluctuations, responsible for finite $v_{2}$ even at zero impact parameter Chatrchyan et al. (2014), while the $b_{n}$ terms reflect the response to global deformation. Analogous relations can be written down for moments of initial conditions:
@@
Both $a_{n}$ and $b_{n}$ (and $a_{n}^{\prime}$ and $b_{n}^{\prime}$) are influenced by how colliding nucleons deposit energy in the overlap region. In phenomenological applications, the energy density is typically parametrized in a flexible way from $T_{\rm A,B}$, with those parameters inferred from experimental measurements. An oftenly used Ansatz is the generalized mean of the Trento model Nijs and van der Schee (2023):
@@
In the simplest scenario of head-on collisions involving spherical nuclei, the initial density distribution closely follows the nucleon density distribution in the $xy$-plane: $T(x,y)\approx T_{A}\approx T_{B}$. Introducing a deformation perturbation along an Euler angle, $T_{A}=T_{A,0}+\delta_{A}(\Omega_{A})$, one can show that the deformation effects are independent of parameter $p$, $T\approx T_{0}+(\delta_{A}(\Omega_{A})+\delta_{B}(\Omega_{B}))/2$. The spherical baseline $T_{0}(x,y)$ determines the $a_{n}$ coefficients, while the deformation-induced perturbations $\delta(x,y)$ vary with the nuclear orientation and control the values of $b_{n}$. In this case, the initial geometry is determined mostly by the nucleon distribution  [^4] However, in non-head-on collisions or when $q\neq 1$, the linear dependence on nuclear deformation does not hold, and the relation between the initial energy distribution and nucleon distribution becomes non-trival.. Additionally, both $a_{n}$ and $b_{n}$ vary with $\sqrt{s_{\mathrm{NN}}}$. Gluon saturation effects at high energy tend to smooth out local fluctuations, reducing $a_{n}$. Conversely, the reduced central density could amplify the influence of nuclear deformation, increasing $b_{n}$.
@@
Implementation of the method. To apply the imaging-by-smashing effectively, we propose an iterative process by utilizing collisions of several isobaric or isobar-like nuclei (see Fig. 5)
@@
  • Prediction. Once the response coefficients are calibrated and validated, we can make discoveries about unknown shapes from the measured flow observables in collisions of nuclei of interest.
@@
This iterative process allows us to establish direct connection between measured flow observables and nuclear shape in Eq. (3), making this method fully data-driven. Within hydrodynamics or transport model framework, one can then systematically improve the implementation initial conditions and explore the alternative definition of $\varepsilon_{n}$ and $d_{\perp}$ to achieve consistent description of the experimental data. We are currently at the calibration step, showing reverse engineering of the state-of-the-art hydro model gives the same shape parameter as low energy STAR Collaboration (2024); Zhang and Jia (2022). More collision species will enable the validation and prediction stages, making the method more towards data-driven.
@@
Figure 5: Implementation of the imaging method. The interative processes of calibration, validation and prediction to established between nuclear structure and final state flow observables or initial conditions structure, as well as tradition approach of constraining the initial condition and QGP properties via hydrodynamic or transport model description of final state flow.
@@
The reverse engineering idea holds at collision energies of a few tens of GeV or higher, where the distinct stages of the collision – initial condition, QGP evolution, and freeze-out – occur on well-separated timescales. At lower energies of below a few GeV, these timescales overlap, the concept of initial condition in terms of $\varepsilon_{n}$ and $d_{\perp}$ are no longer valid. Instead nuclear structure themselves become the initial conditions of the collisions. Despite these complexities, Eq. (3) still remain valid, except that the response coefficients $a_{n}$ and $b_{n}$ need to be calibrated using collisions of isobar species at the same $\sqrt{s_{\mathrm{NN}}}$. The ability to vary the nuclear shape, measure the flow response, and compare with model prediction remain a valuable approach for investigating the dynamics and properties of nuclear matter at lower energies. For instance, deformed nuclei offer a unique way to globally rearrange nucleons, allowing collisions of such nuclei to probe the nuclear equation of state. The Fermi momenta of nucleons are also relevant to the measured observables, leading to a more intricate interplay between nuclear structure and collision dynamics.
Imaging-by-smashing as a discovery tool. The imaging-by-smashing technique offers exciting potential for advancing both heavy-ion and nuclear structure research Jia et al. (2024). Realizing its full potential, however, requires rigorous calibration and a thorough understanding of the QGP’s initial condition and evolution – objectives already central to heavy-ion physics. By systematically studying a selected set of isobaric or isobar-like nuclei with well-known structural properties, we gain crucial “lever arms” to calibrate and validate this method. Once calibrated, imaging-by-smashing can be a powerful tool for exploring the ground-state structure of atomic nuclei. By colliding species whose shape parameters are difficult to determine through low-energy methods and analyzing the resulting heavy-ion observables, nuclear shapes can be directly extracted. The large number of charged particles produced in each event facilitates the study of higher-order nuclear deformations, including octupole and hexadecapole shapes, through measurements of higher-order flow harmonics. Moreover, measurements of multi-particle correlation enables a sepration between static deformation from dynamical shape fluctuations.
@@
One exciting application involves “shape coexistence”, the phenomenon where some nuclei have multiple distinct shapes with nearly degenerate ground-state ($0^{+}$) energies Garrett et al. (2022). The nuclear ground-state wavefunction becomes an admixture of these different shape configurations. For instance, a superposition of two shapes can be represented as:
@@
where the mixing angle $\eta$ can be determined from electric quadrupole and monopole transitions between the two shape eigenstates. Some stable nuclei, such as 42Ca, 74Kr, and 152Sm, exhibit very strong mixing of two distinct shapes in their ground states, representing unique stable quantum entanglement of collective many-body wavefunctions. High-energy collisions effectively “projects” nuclei onto one specific, instantaneous shapes. It’s not that the nucleus becomes that shape; rather, the collision reveals the particular shape configuration that was present at the moment of impact. The presence of shape coexistence, depending on how this projection process is realized, could lead to different expectations in the measured flow fluctuations. High-energy collisions provides a unique setting to probe quantum entanglement of collective many-body wavefunctions.

### v4

@@ Abstract
High-energy nuclear collisions have recently emerged as a promising “imaging-by-smashing” approach to reveal the intrinsic shapes of atomic nuclei. Here, I outline a conceptual framework for this technique, explaining how nuclear shapes are encoded during quark-gluon plasma formation and evolution, and how they can be decoded from final-state particle distributions. I highlight the method’s potential to advance our understanding of both nuclear structure and quark-gluon plasma physics.
@@
Imaging techniques are fundamental to scientific discovery, enabling us to visualize and understand structures across vastly different scales. These methods employ external probes, often photons, to interact with samples and glean structural information. The sample’s structural information is encoded in the scattered probe. In X-ray crystallography, for example, a molecule’s three-dimensional organization is deduced from momentum-space scattering patterns. The imaging process takes place as the probe traverses the sample, and because any structural disintegration happens only after scattering, the captured information remains intact.
@@
A radically different approach has emerged in nuclear physics: rather than using external probes, atomic nuclei are collided at ultra-high energies to create a hot, dense state of matter called the quark-gluon plasma (QGP). Remarkably, the geometric structure of the colliding nuclei determines the QGP’s initial conditions and influences its subsequent evolution. This process ultimately imprints specific patterns on the momentum distributions of thousands of final-state particles. By carefully “rewinding” this evolution, one can reconstruct an effective image of the original nuclei.
@@
The STAR Collaboration demonstrated this concept by extracting the shape of 238U from 238U+238U collision data that are consistent with low-energy measurements. Despite completely destroying the nuclei, their structural information is preserved with sufficient detail to allow reconstruction STAR Collaboration (2024). This breakthrough raises fundamental questions: How are nuclear shapes encoded in such violent collisions? How can they be deduced from final-state particles? What are the broader implications for nuclear physics?
@@
Traditional methods for nuclear shapes. Atomic nuclei are bound systems of protons and neutrons (nucleons) held together by the strong nuclear force. While often depicted as spheres, nuclei can adopt prolate, oblate, or even pear-like configurations – shapes that reflect their underlying many-body wavefunctions. These wavefunctions govern the positions and momenta of nucleons. The non-perturbative nature of the strong force makes theoretical predictions of these shapes challenging and experimental input crucial.
@@
Unlike molecules, nuclei cannot be oriented in a crystalline lattice for direct coherent imaging. Instead, low-energy techniques such as electron scattering, laser spectroscopy, and Coulomb excitation (Coulex) have been used to infer nuclear shapes Yang et al. (2023). Electron scattering probes nuclei one at a time. This provides an orientation-averaged image where nuclear deformation appears as modifications to charge distributions. Laser spectroscopy probes nuclear shapes through measurements of hyperfine atomic transitions. Coulex uses low-energy ions to excite nuclei into rotational and vibrational states. Nuclear shapes are then deduced by detecting gamma rays emitted during relaxation and comparing the measured emission rates to theoretical models.
@@
However, these methods rely on electromagnetic interactions and thus cannot directly probe neutron distributions. Moreover, their relatively long timescales yield time-averaged, blurred representations of nuclear shapes, akin to long-exposure photographs. This temporal averaging obscures rapid fluctuations and dynamical variations in nucleon distributions that affect nuclear shapes.
@@
From a quantum mechanical perspective, a nucleus in its isolated ground state exists in a superposition of all possible orientations, making its intrinsic shape unobservable Verney (2025). However, high-energy heavy-ion collisions represent strongly interacting, dynamical systems where nuclear ground states are not eigenstates of the combined system. The extremely rapid collision process acts as a quantum measurement, effectively collapsing the nuclear many-body wavefunction and projecting nucleons into specific positions, from which nuclear shape and orientation can be defined.
Initially, the nucleus’s shape directly influences the conditions under which the QGP forms. As the plasma expands hydrodynamically, information about the original nuclear shape transforms and becomes encoded in the momentum distribution of thousands of final-state particles. By analyzing these distributions and reverse-engineering the expansion process event by event, nuclear shapes can be reconstructed. The precision of this method hinges on how well the initial conditions, equation of state, and transport properties of the QGP are understood. These aspects have been the focus of the heavy-ion community for decades and are now believed to be well-constrained Busza et al. (2018).
@@
Despite its seemingly extreme nature, high-energy nuclear imaging may offer distinct advantages. First, the collision probes the ground state nuclear many-body wavefunction in position space, information not readily accessible through more conventional means. The colliding nuclei and their Coulomb fields are Lorentz-contracted, ensuring that the nuclei are causally disconnected before the collision. This causal disconnection prevents pre-collision excitation effects that could contaminate the measurement. While nuclei could be Coulomb excited when they pass by each other, most isomeric states are very short-lived (typically $10^{-12}$ s) and cannot persist in the circulating beam long enough to participate in the subsequent collisions  [^1] Some isomers have exceptionally long lifetime, such as 7.7s 11/2- state of 197Au and 8.9 days 11/2- state of 129Xe, which can accumulate in the beam. However, no estimates are currently available on their production rates in high-energy accelerators.
@@
Energy dependence of the nuclear image. The observed nuclear image, and its effective many-body wavefunction, depends critically on the collision energy $\sqrt{s_{\mathrm{NN}}}$, which determines the effective “shutter speed” of the imaging process. At low energies, where the shutter speed is slow, collective rotational and vibrational DOF dominate the nucleus’s apparent shape. For instance, the rotational DOF of a deformed heavy nucleus, with excitation energies around 0.1 MeV, corresponds to timescales of $\tau\sim 10^{3}-10^{4}$ fm/$c$. As collision energy increases, faster modes such as nucleon clustering and short-range correlations come into play. At still higher energies, the resonance structure of nucleons emerges, eventually giving way to the subnucleonic quarks and gluons. The imaging process effectively captures all DOF slower than the nuclear crossing timescale. As a result, varying $\sqrt{s_{\mathrm{NN}}}$ provides a natural way to study the evolution of the nuclear wavefunction across energy scales.
@@
While electron-nucleus scattering probes the one-body distribution of the scattering centers – protons at low energy and quarks and gluons at high energy – much less is known about the nuclear many-body distributions across energy scales. High-energy heavy-ion collisions help fill this gap. Measurements of anisotropic flow coefficients reveal quantum fluctuations associated with the finite number of scattering centers Heinz (2013). These quantum fluctuations induce deformation in the transverse ($xy$) nucleon distribution, with eccentricities scaling roughly as $1/\sqrt{A}$, where $A$ is the mass number  [^2] The eccentricities are mostly determined by the positions of nucleons, as most quarks and gluons are confined within them and nuclear modifications are modest.. This fluctuation-driven deformation produces non-zero flow even in head-on collisions Chatrchyan et al. (2014), despite the nuclear overlap region being isotropic on average (see Fig. 2).
@@
Figure 3: Relation between initial- and final-state of high-energy nuclear collisions. The features of collision geometry are characterized by its shape and size parameters, $\mathcal{E}_{n}$ and $d_{\perp}$ ($\bf a$}). They are linearly related to observables that describe the transverse momentum $p_{\mathrm{T}}$ spectra in each event, i.e., the anisotropic flow coefficients $V_{n}$ and the average transverse momentum $p_{\mathrm{T}}]$ ($\bf b$}). The event-to-event variations of these initial- and final-state quantities are linearly related.
Reverse-engineering nuclear shape. Having established the conceptual foundation, we now turn to the practical implementation of nuclear shape reconstruction. A nucleus with quadrupole deformation can be described by a surface function in terms of the polar angle $\theta$ and azimuthal angle $\phi$,
@@
Figure 4: Three steps in the imaging-by-smashing method. Illustrated here are collisions of spherical nuclei ($\bf a1$$\bf a3$}) and prolate-deformed nuclei with $\beta_{2}=0.28$ ($\bf b1$$\bf b3$}) for two representative collision events: ($\bf a1$} and $\bf b1$}) initial configurations of the colliding nuclei, ($\bf a2$} and $\bf b2$}) initial geometry in the transverse plane, and ($\bf a3$} and $\bf b3$}) final-state distribution of particles in azimuthal angle and $p_{\mathrm{T}}$. The event-to-event variation in the initial-state geometry and final-state distributions is quantified by the moments involving $\varepsilon_{n}$ and $d_{\perp}$ of the initial state and $v_{n}$ and $p_{\mathrm{T}}]$ of the final state, as indicated in the middle row. Parameter $a$ represents the nuclear surface diffusivity. The nucleon positions are simulated in a Monte-Carlo Glauber model with $R_{0}=6.81$ fm, $a=0.55$ fm, and $A=238$. Imaging involves reconstructing the initial geometry from final-state particles (Step 1), relating it to the nuclear shape, which is affected by both global deformation and quantum fluctuations (Step-2), and comparing different collision systems to isolate the global deformation (Step-3).
Step1: From the final state to the initial condition. The QGP’s initial geometry in a single event is characterized by its energy density distribution $T(x,y)$ (Fig. 3a), where $x$ and $y$ are distribution’s principal axes relative to its center-of-mass. Key parameters include the total energy $E=\int T(x,y)dxdy$, the elliptic and triangular eccentricities $\varepsilon_{2}$ and $\varepsilon_{3}$, and inverse of mean-square area $d_{\perp}$:
@@
with $\left[..\right]$” indicates averaging weighted by $T(x,y)$, and $\Phi_{n}$ represents the orientation of the $n^{\rm th}$ eccentricity.
The total energy $E$ influences the number of charged particles. Meanwhile, the initial eccentricites $\varepsilon_{n}$ drive anisotropic flow, described by $dN/d\phi~\propto~1+2\sum_{n}v_{n}\cos(n(\phi-\Psi_{n}))$, where $v_{n}$ (with phase $\Psi_{n}$) define the elliptic ($v_{2}$) and triangular ($v_{3}$) flow coefficients (Fig. 3b). Additionally, $d_{\perp}$ governs the radial expansion or “radial flow”, affecting the average transverse momentum $p_{\mathrm{T}}]$ of final-state particles. Hydrodynamic modeling suggests approximately linear relationships between these initial- and final-state observables. Specifically, $v_{n}\propto\varepsilon_{n}$ and $\delta p_{\mathrm{T}}/p_{\mathrm{T}}\propto\delta d_{\perp}/d_{\perp}$ Niemi et al. (2013); Boźek and Broniowski (2012)  [^3] Here, $\frac{\delta p_{\mathrm{T}}}{p_{\mathrm{T}}}=\frac{[p_{\mathrm{T}}]-\left\langle[p_{\mathrm{T}}]\right\rangle}{\left\langle[p_{\mathrm{T}}]\right\rangle}$ and $\frac{\delta d_{\perp}}{d_{\perp}}=\frac{d_{\perp}-\left\langle d_{\perp}\right\rangle}{\left\langle d_{\perp}\right\rangle}$ denote event-by-event fluctuations, and $\left\langle..\right\rangle$” indicates average over events.. The proportionality constants (response coefficients) are strongly impacted by QGP properties.
Event-by-event variations in nuclear geometry arise from both global deformation and quantum fluctuations. These variations are quantified by moments of initial- and final-state observables. These moments are connected by the linear response relations shown in Fig. 3. Our understanding of the initial conditions, evolution, and properties of the QGP is mainly derived from the study of these relations Heinz (2013); Busza et al. (2018). Many such moments can be used, as indicated in the middle row of Fig. 4. The STAR collaboration specifically focused on three: $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ STAR Collaboration (2024).
@@
Step2: from the initial conditions to nuclear shape. The QGP’s initial condition is closely tied to nucleon distributions in colliding nuclei A and B, described by thickness functions $T_{\rm A,B}(x,y)$, which fluctuate from event to event. In the presence of global deformation, these functions also depend on the nuclei’s orientations prior to collision (see Fig. 4a1 and 4b1). For head-on collisions, random orientations of prolate deformed nuclei lead to significant, anti-correlated fluctuations between $v_{2}$ and $p_{\mathrm{T}}]$ Giacalone (2020). General considerations imply that such fluctuations follow simple parametric forms Jia (2022b). Specifically, the three moments used by the STAR Collaboration are  [^4] Shape fluctuations can be accounted for by replacing $\beta_{2}^{2}$ and $\beta_{2}^{3}\cos(3\gamma)$ by $\left\langle\beta_{2}^{2}\right\rangle$ and $\left\langle\beta_{2}^{3}\cos(3\gamma)\right\rangle$, respectively.:
@@
Here, $a_{n}$ and $b_{n}$ are positive coefficients that depend on the impact parameter. The $a_{n}$ terms capture contributions from quantum fluctuations, responsible for finite $v_{2}$ even at zero impact parameter Chatrchyan et al. (2014), while the $b_{n}$ terms reflect the response to global deformation. Analogous relations can be written down for moments of initial conditions:
@@
Both $a_{n}$ and $b_{n}$ (and $a_{n}^{\prime}$ and $b_{n}^{\prime}$) are influenced by how colliding nucleons deposit energy in the overlap region. In phenomenological applications, the energy density is typically parametrized in a flexible way from $T_{\rm A,B}$, with those parameters inferred from experimental measurements. A commonly used ansatz is the generalized mean of the Trento model Nijs and van der Schee (2023):
@@
Consider the simplest scenario of head-on collisions involving spherical nuclei. In this case, the initial density distribution closely follows the nucleon density distribution in the $xy$-plane: $T(x,y)\approx T_{A}\approx T_{B}$. Introducing a deformation perturbation along an Euler angle, $T_{A}=T_{A,0}+\delta_{A}(\Omega_{A})$, one can show that the deformation effects are independent of parameter $p$, $T\approx T_{0}+(\delta_{A}(\Omega_{A})+\delta_{B}(\Omega_{B}))/2$. The spherical baseline $T_{0}(x,y)$ determines the $a_{n}$ coefficients. Meanwhile, the deformation-induced perturbations $\delta(x,y)$ vary with nuclear orientation and control the values of $b_{n}$. In this case, the initial geometry is determined mostly by the nucleon distribution  [^5] However, in non-head-on collisions or when $q\neq 1$, the linear dependence on nuclear deformation does not hold, and the relation between the initial energy distribution and nucleon distribution becomes non-trivial.. Additionally, both $a_{n}$ and $b_{n}$ vary with $\sqrt{s_{\mathrm{NN}}}$. Gluon saturation effects at high energy tend to smooth out local fluctuations, reducing $a_{n}$. Conversely, the reduced central density could amplify the influence of nuclear deformation, increasing $b_{n}$.
@@
Implementation of the method. To develop the imaging-by-smashing method toward practical application, we propose an iterative process by utilizing collisions of several isobaric or isobar-like nuclei (see Fig. 5)
@@
  • Prediction. Once the response coefficients are calibrated and validated, we may potentially make discoveries about unknown shapes from the measured flow observables in collisions of nuclei of interest.
@@
This iterative process aims to establish a direct connection between measured flow observables and nuclear shape in Eq. (3), making this method fully data-driven. Within hydrodynamic or transport model frameworks, one can then systematically improve the implementation of initial conditions and explore the alternative definition of $\varepsilon_{n}$ and $d_{\perp}$ to achieve a consistent description of the experimental data. We are currently at the early calibration step, where reverse engineering using state-of-the-art hydrodynamic models shows encouraging agreement with shape parameters from low-energy methods STAR Collaboration (2024); Zhang and Jia (2022). More collision species will enable the validation and prediction stages, making the method more data-driven.
@@
Figure 5: Implementation of the imaging method. The iterative processes of calibration, validation, and prediction to establish connections between nuclear structure and final-state flow observables or initial conditions structure, as well as the traditional approach of constraining initial conditions and QGP properties via hydrodynamic or transport model description of final state flow.
@@
The reverse engineering idea holds at collision energies of a few tens of GeV or higher, where the distinct stages of the collision – initial condition, QGP evolution, and freeze-out – occur on well-separated timescales. At lower energies of below a few GeV, these timescales overlap, and the concept of initial conditions in terms of $\varepsilon_{n}$ and $d_{\perp}$ is no longer valid. Instead, nuclear structures themselves become the initial conditions of the collisions. Despite these complexities, Eq. (3) remains valid, except that the response coefficients $a_{n}$ and $b_{n}$ need to be calibrated using collisions of isobar species at the same $\sqrt{s_{\mathrm{NN}}}$. The ability to vary the nuclear shape, measure the flow response, and compare with model predictions remains a valuable approach for investigating the dynamics and properties of nuclear matter at lower energies. For instance, deformed nuclei offer a unique way to globally rearrange nucleons, allowing collisions of such nuclei to probe the nuclear equation of state. The Fermi momenta of nucleons are also relevant to the measured observables, leading to a more intricate interplay between nuclear structure and collision dynamics.
Imaging-by-smashing as a potential discovery tool. The imaging-by-smashing technique offers exciting potential for advancing both heavy-ion and nuclear structure research Jia et al. (2024). However, realizing this potential requires rigorous calibration and thorough understanding of QGP initial conditions and evolution – objectives already central to heavy-ion physics. A systematic study of isobaric or isobar-like nuclei with well-known structural properties shall provide crucial calibration benchmarks for validating this method.
It is important to clarify what this method measures. Rather than extracting intrinsic nuclear shapes via non-invasive methods as in low-energy nuclear physics, this technique extracts effective shape parameters that reflect how the underlying many-body wavefunction influences the outcomes of the dynamical collision process. The nuclear structures accessed at high-energy heavy ion collisions are not entirely the same as those at low energies. While inherently model-dependent, I am hopeful that systematic validation across multiple collision systems will establish rigorous connections between nuclear structure and collision dynamics, making it a valuable complement to low-energy measurements.
Once properly calibrated and validated, imaging-by-smashing can become a powerful tool for exploring nuclear structure. The method is valuable for probing species whose shape parameters are difficult to determine through conventional techniques. The large number of particles produced per event enables exploration of higher-order deformations, including octupole and hexadecapole shapes, through higher-order flow harmonics. Multi-particle correlations can also be used to separate static deformation from dynamical shape fluctuations.
@@
A particularly exciting application involves “shape coexistence”, the phenomenon where some nuclei have multiple distinct shapes with nearly degenerate ground-state ($0^{+}$) energies Garrett et al. (2022). The nuclear ground-state wavefunction becomes an admixture of these different shape configurations. For instance, a superposition of two shapes can be represented as:
@@
where the mixing angle $\eta$ can be determined from electric quadrupole and monopole transitions between the two shape eigenstates. Some stable nuclei, such as 42Ca, 74Kr, and 152Sm, exhibit very strong mixing of two distinct shapes in their ground states. This represents a unique stable quantum superposition of collective many-body wavefunctions. Each high-energy collision should effectively collapse the nuclear wavefunction onto one specific shape component. The collision doesn’t transform the nucleus into that shape; rather, it reveals the particular shape configuration present at the moment of impact. The presence of shape coexistence, depending on how this projection process is realized, could lead to different expectations in the measured flow fluctuations. High-energy collisions provide a unique setting to probe quantum superposition of collective wavefunctions, though this capability requires further theoretical and experimental development.